Erin is a 28-year-old professional woman who is 5 feet 8 inches tall and maintains her weight at 118 pounds by following a lacto-ovo (non-fat milk and egg whites only) vegetarian diet that supplies approximately 1200 calories a day. With her understanding that protein should provide between 10 and 35 percent of her daily calories, she reasons that her daily intake of 40 grams of protein from milk, eggs, legumes, and nuts is adequate for her needs. She is concerned, however, that she has been sick more than usual and has experienced two stress fractures in her leg over the past three years while exercising. For questions 1, 2, 3, and 7 – please show your work/calculations.

1. What is the protein RDA per kilogram of healthy body weight? a. 0.4 g b. 0.6 g c. 0.8 g d. 1 g e. 1.2 g

2. Assuming a healthy weight for Erin is 141 pounds, what is her recommended daily protein requirement? a. 41 g b. 46 g c. 51 g d. 56 g e. 61 g

3. Remembering that protein provides 4 kcalories per gram, what percentage of Erin’s current total energy intake comes from protein? a. 7.3% b. 9.3% c. 11.3% d. 13.3% e. 15.3%

4. What percentage of total energy intake should come from protein?
Answer #1

Answer 1:-

For a healthy body weight, the recommended RDA per kilogram is 0.8 gms. This is a universally recommended measure.

So, correct option is C.

Answer 2:-

Calculations:-

Body weight = 141 lbs

Converting pounds to kilogram we get 141/2.2 = 64.09 kgs

RDA per kg of body weight is 0.8 gms/kg

Recommended daily protein intake = Weight (kgs) × RDA for protein

i.e = 64.09 × 0.8 = 51.27 gms of protein which overall counts to 51gms.

So, correct option is option C (51gms)

Answer 3:-

According to me, it should be 1200 kcal to be mentioned. If she is supplied by only 2000 calculated it means that it would be around 2 kcal per day which will not meet the needs.

She consumes 40 gms of protein and each gram provided 4 kcal. This, 40 gms of protein will provide around 160 kcal of energy.

Total energy input is 1200 kcal. So out of it, 160 kcal is provided by protein which accounts to 0.133 and multiplied by 100 the value comes to 13.3%.

Energy intake from protein = 160/1200 × 100 = 13.3 %

So , correct option is option D.